Solve for x: (x - 4)/2 - 1/5 = (7x + 1)/20
Combine like terms: 3x - 6x^2 + 2 - 4x^2 + 7x - 1 - 8 - 3x^2
Name the property of rreal numbers illustrated below:
4 + (7 + 1) = (4 + 7) + 1

Answer by sdmmadam@yahoo.com(530)

1)Solve for x: (x - 4)/2 - 1/5 = (7x + 1)/20
(x - 4)/2 - 1/5 = (7x + 1)/20
Multiplying by 20 which is the lcm of 2,5 and 20
10(x-4)-4X(1)=(7x+1)
10x-40-4=7x+1
10x-44=7x+1
10x-7x=1+44 (grouping like terms, here side changed and hence sign changed)
3x=45
x=45/3=15
Verification: putting x=15 in the given equation,we get LHS = (55-2)/10 = 53/10
and RHS = (7X15+1)/20 = (105+1)/20 = 106/20 = 53/6
2)Combine like terms: 3x - 6x^2 + 2 - 4x^2 + 7x - 1 - 8 - 3x^23x - 6x^2 + 2 - 4x^2 + 7x - 1 - 8 - 3x^2
=(-6x^2-4x^2-3x^2)+(3x+7x)+(2-1-8) (three +two+three terms = 8 terms in all)
=(-6-4-3)X^2+(3+7)x+(2-9)
=-13x^2+10x-7
[Can we pluck any term from anywhere in the given expression and place it near any other term we wish ?
YES. That is the advantage of additive commutativity and additive associativity.. By commutativity applied repeatedly we bring two terms of like powers of x to gether.
You may visualise friends visiting a cinema hall seated in a row and doing this act of slowly moving place two friends in action at a time swapping the immediate places and this act done repeatedly will fetch the desired result of a friend getting seated next to a desired friend.]
3)Name the property of rreal numbers illustrated below:
4 + (7 + 1) = (4 + 7) + 1
This property: A+(B+C) = (A+B) +C
is called the Associativity property with respect to addition
and it holds good in the set of real number